[Brahms-l] Shift report 20060611 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 11 2006 - 16:09:17 EDT
Shifter leader was kh
Shifters were nk


Had beam during most of the shift.
We ran FS at 3 deg 1/2A and MRS at 45 deg 4kG B.
T3 Anodes tripped once as noted in elog.
As indicated in last shift, t3 A1-xu is reading 5249 V.  We strongly suspect a readout error on the mainframe.  Everything else seems OK with t3.
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 11 16:09:45 2006

This archive was generated by hypermail 2.1.8 : Sun Jun 11 2006 - 16:09:59 EDT